OpenStudy (anonymous):

Connections Academy help please? 2. Use the Change of Base Formula to evaluate log[8]77. 3.467 1.886 4.344 2.089

OpenStudy (anonymous):

\[\log_8(77)=\frac{\ln(77)}{\ln(8)}\] and a calculator

OpenStudy (anonymous):

I don't have my calculator. Can you tell me the answer?

OpenStudy (anonymous):

Unit 7 Lesson 6!!

OpenStudy (zehanz):

Because you do not really have to calculate anything, you can make an educated guess: \[\log_{8}77=a <=> 8^a=77 \] Now 8^2 =64, so 8^2.... = 77, because 77 is only slightly larger than 64. The only answer that fits is therefore 2.089.

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

Yes, I'm sure!

OpenStudy (zehanz):

8 = 8^1, so log(8)8=1 64=8^2, so log(8)64=2 512=8^3, so log(8)512=3 See? That's why log(8)77 must be only a little more than 2!